# Data-driven decision-making under uncertainty with entropic risk measure

### v1

@@ A.2 Proof of Theorem 1
We want to show that $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))+\delta_{N}(\mathbb{Q}_{N})\rightarrow\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))$ almost surely. To do so, we will show that both $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))\rightarrow\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))$ and $\delta_{N}(\mathbb{Q}_{N})\rightarrow 0$ almost surely. Indeed, if both $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))\rightarrow\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))$ and $\delta_{N}(\mathbb{Q}_{N})\rightarrow 0$ almost surely, then we can use the fact that the sum of two convergent sequence converge to the sum of their limits to conclude that:
@@
Step 1: $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))\rightarrow\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))$ almost surely. Given that ${\mathbb{E}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]$ is finite (see Lemma 1) and each $\xi_{i}$ is i.i.d., the strong law of large numbers tells us that
@@
Since $\log$ is continuous over the strictly positive values, using the continuous mapping theorem (Van der Vaart,, 2000), we obtain:
(22)$\Rightarrow\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))\right)\rightarrow\frac{1}{\alpha}\log\left({\mathbb{E}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]\right)\text{ almost surely}.$
@@
Let’s consider any sequence {¯$\{\bar{\mathbb{Q}}_{N}\}_{N=1}^{\infty}$ that is uniformly exponentially bounded, with each element of the sequence satisfying Assumption 2. For each $N$, we have
$\exp(-\alpha\delta_{N}(\bar{\mathbb{Q}}_{N}))$ $\ =\ \exp(\alpha\mbox{median}[\rho_{\hat{\bar{\mathbb{Q}}}_{N,N}}(\zeta)-\rho_{\bar{\mathbb{Q}}_{N}}(\zeta)|\bar{\mathbb{Q}}_{N}])$
$\ =\ \mbox{median}[\exp(\alpha\rho_{\hat{\bar{\mathbb{Q}}}_{N,N}}(\zeta)-\alpha\rho_{\bar{\mathbb{Q}}_{N}}(\zeta))|\bar{\mathbb{Q}}_{N}]$
$\ =\ \mbox{median}\left[\exp\left(\log\left(\frac{(1/N)\sum_{i=1}^{N}\exp(\alpha\zeta_{i})}{{\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]}\right)\right)\;\big{|}\;\bar{\mathbb{Q}}_{N}\right]$
@@
where $X_{N}:=\frac{(1/N)\sum_{i=1}^{N}\exp(\alpha\zeta_{i})}{{\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]}$. The second equality follows by the monotonicity of the exponential function, and the third equality follows by the definition of the entropic risk and the properties of the logarithm. Note that $\{\zeta_{i}\}_{i=1}^{N}$ are drawn i.i.d. from ¯$\bar{\mathbb{Q}}_{N}$.
To analyze the $\mbox{median}[X_{N}|\bar{\mathbb{Q}}_{N}]$, we first study the mean and variance of $X_{N}$. It is easy to see that ${\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[X_{N}]=1$, while the variance of $X_{N}$ can be bounded as follows,
$\begin{array}[]{ll}\mbox{Var}_{\bar{\mathbb{Q}}_{N}}[X_{N}]&=\displaystyle\frac{\mbox{Var}_{\bar{\mathbb{Q}}_{N}}[\sum_{i=1}^{N}\exp(\alpha\zeta_{i})]}{\left(N{\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]\right)^{2}}\\[12.91663pt] &=\displaystyle\frac{\mbox{Var}_{\bar{\mathbb{Q}}_{N}}[(\exp(\alpha\zeta)]}{N\left({\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]\right)^{2}}\\[12.91663pt] &\leq\displaystyle\frac{2G\exp(\frac{4G}{C})}{N(C-2)},\end{array}$
where the second equality follows from the fact that the $\{\zeta_{i}\}_{i=1}^{N}$ are i.i.d. The inequality follows since {¯$\{\bar{\mathbb{Q}}_{N}\}_{N=1}^{\infty}$ satisfy Assumption 2, thus Lemma 1 provides bounds for both $\mbox{Var}_{\bar{\mathbb{Q}}_{N}}[(\exp(\alpha\zeta)]$ and ${\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]$, resulting in the bound $\frac{\text{Var}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]}{[{\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]]^{2}}\leq\frac{2G\exp(\frac{4G}{C})}{C-2}$.
We next show that $\mbox{median}[X_{N}|\bar{\mathbb{Q}}_{N}]$ is bounded. To this end, consider the Chebyshev’s inequality
¯${\mathbb{P}}_{\bar{\mathbb{Q}}_{N}}\left(|X_{N}-1|\geq k\sqrt{\mbox{Var}_{\bar{\mathbb{Q}}_{N}}[X_{N}]}\right)\leq\frac{1}{k^{2}}.$
Substituting the bound for the $\mbox{Var}_{\bar{\mathbb{Q}}_{N}}[X_{N}]$, and setting $k=2$ which implies an upper tail probability bound of 25%, results in
@@
We conclude from the above analysis that any sequence {¯$\{\bar{\mathbb{Q}}_{N}\}_{N=1}^{\infty}$ that is uniformly exponentially bounded as prescribed in Assumption 2 must be such that $\exp(-\delta_{N}(\bar{\mathbb{Q}}_{N}))=\mbox{median}[X_{N}|\bar{\mathbb{Q}}_{N}]$ converges to 1 almost surely. Since Assumption 2 imposes that such sequences are almost surely obtained, we must have that $\delta_{N}(\mathbb{Q}_{N})\rightarrow 0$ almost surely.
A.3 Proof of Theorem 2
(15) is equivalent to
$\mathop{\rm min}_{{{\boldsymbol{t}}}}\quad$ $\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i\in[N]}\exp(\alpha t_{i})\right)$ (23a)
subject ${\rm subject\mbox{ }\rm to}\quad$ $t_{i}\geq\ell({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i})$ $\forall\hat{{{\boldsymbol{\xi}}}}_{i}\in\Xi^{i}_{\epsilon},\;\forall i\in[N],$ (23b)
where $\Xi^{i}_{\epsilon}=\{{{\boldsymbol{\xi}}}:\|{{\boldsymbol{\xi}}}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon\}$. Since $\ell({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i})=\mathop{\rm max}_{j\in[m]}\ell_{j}({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i})$, we obtain:
$\mathop{\rm min}_{{{\boldsymbol{t}}}}\quad$ $\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i\in[N]}\exp(\alpha t_{i})\right)$ (24a)
subject ${\rm subject\mbox{ }\rm to}\quad$ $t_{i}\geq\ell_{j}({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i})$ $\forall\hat{{{\boldsymbol{\xi}}}}_{i}\in\Xi^{i}_{\epsilon},\;\forall i\in[N],j\in[m].$ (24b)
The relative interior of the intersection of set $\Xi^{i}_{\epsilon}$ and domain of $\ell_{j}({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i})$ is non-empty for all $\epsilon\geq 0$. So, we can use Fenchel duality theorem (Ben-Tal et al.,, 2015) to obtain
$\mathop{\rm sup}_{\hat{{{\boldsymbol{\xi}}}}_{i}\in\Xi_{\epsilon}^{i}}\ell_{j}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})=\mathop{\rm inf}_{{{\boldsymbol{\varphi}}}_{ij}}\delta^{*}({{\boldsymbol{\varphi}}}_{ij}\rvert\Xi_{\epsilon}^{i})-\ell_{j*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}_{ij})$ (25)
where ${{\boldsymbol{\varphi}}}_{ij}\in\mathbb{R}^{d}$, $\ell_{j*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}_{ij}):=\mathop{\rm inf}_{{{\boldsymbol{\xi}}}}{{\boldsymbol{\varphi}}}_{ij}^{\top}{{\boldsymbol{\xi}}}-\ell_{j}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is the partial concave conjugate of $\ell_{j}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ and $\delta^{*}({{\boldsymbol{\varphi}}}_{ij}\rvert\Xi^{i}_{\epsilon})$ is the support function of $\Xi_{\epsilon}^{i}$, i.e.,
$\delta^{*}({{\boldsymbol{\varphi}}}_{ij}\rvert\Xi^{i}_{\epsilon})=\mathop{\rm sup}_{{{\boldsymbol{\xi}}}\in\Xi^{i}_{\epsilon}}{{\boldsymbol{\varphi}}}_{ij}^{\top}{{\boldsymbol{\xi}}}={{\boldsymbol{\varphi}}}_{ij}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i}+\mathop{\rm sup}_{{{\boldsymbol{\Delta}}}:\|{{\boldsymbol{\Delta}}}\|\leq\epsilon}{{\boldsymbol{\varphi}}}_{ij}^{\top}{{\boldsymbol{\Delta}}}={{\boldsymbol{\varphi}}}_{ij}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i}+\epsilon\|{{\boldsymbol{\varphi}}}_{ij}\|_{*}$ (26)
by the definition of the dual norm. We therefore obtain:
$\begin{array}[]{ll}\text{\rm minimize}&\displaystyle\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha t_{i})\right)\\ {\rm subject\mbox{ }\rm to}&\boldsymbol{z}\in\mathcal{Z},\,\boldsymbol{\varphi}_{ij}\in\mathbb{R}^{d}\quad\forall i\in[N],\,j\in[m]\\ &\displaystyle\hat{{{\boldsymbol{\xi}}}}_{i}^{\top}{{\boldsymbol{\varphi}}}_{ij}+\epsilon\|{{\boldsymbol{\varphi}}}_{ij}\|_{*}-\ell_{j*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}_{ij})\leq t_{i}\,,\;\forall i\in[N],\,j\in[m].\end{array}$ (27)

### v3

@@ A.2 Proof of Theorem 2
Proof. We want to show that $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))+\delta_{N}(\mathbb{Q}_{N})\rightarrow\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))$ almost surely. To do so, we will show that both $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))\rightarrow\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))$ and $\delta_{N}(\mathbb{Q}_{N})\rightarrow 0$ almost surely. Indeed, if both $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))\rightarrow\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))$ and $\delta_{N}(\mathbb{Q}_{N})\rightarrow 0$ almost surely, then we can use the fact that the sum of two convergent sequences converges to the sum of their limits to conclude that:
@@
Step 1: $\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))\rightarrow\rho_{\mathbb{P}}(\ell({{\boldsymbol{\xi}}}))$ almost surely. Given that ${\mathbb{E}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]$ is finite (see Lemma 1) and each ${{\boldsymbol{\xi}}}_{i}$ is i.i.d., the strong law of large numbers tells us that:
@@
Since logarithm function is continuous over the strictly positive values, using the continuous mapping theorem (Van der Vaart,, 2000), we obtain:
(21)$\implies\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))\right)\rightarrow\frac{1}{\alpha}\log\left({\mathbb{E}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]\right)\text{ almost surely}.$
@@
Let’s consider any sequence {¯$\{\bar{\mathbb{Q}}_{N}\}_{N=1}^{\infty}$ that is uniformly exponentially bounded, with each element of the sequence satisfying Assumption 2. For each $N$, we have:
$\exp(-\alpha\delta_{N}(\bar{\mathbb{Q}}_{N}))$ $\ =\ \exp(\alpha\,\mbox{median}(\rho_{\hat{\bar{\mathbb{Q}}}_{N,N}}(\zeta)-\rho_{\bar{\mathbb{Q}}_{N}}(\zeta)\lvert\bar{\mathbb{Q}}_{N}))$
$\ =\ \mbox{median}(\exp(\alpha\rho_{\hat{\bar{\mathbb{Q}}}_{N,N}}(\zeta)-\alpha\rho_{\bar{\mathbb{Q}}_{N}}(\zeta))|\bar{\mathbb{Q}}_{N})$
$\ =\ \mbox{median}\left(\exp\left(\log\left(\frac{(1/N)\sum_{i=1}^{N}\exp(\alpha\zeta_{i})}{{\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]}\right)\right)\middle|\bar{\mathbb{Q}}_{N}\right)$
@@
where $X_{N}:=\frac{(1/N)\sum_{i=1}^{N}\exp(\alpha\zeta_{i})}{{\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]}$. The second equality follows by the monotonicity of the exponential function, and the third equality follows by the definition of the entropic risk and the properties of the logarithm function. Note that $\{\zeta_{i}\}_{i=1}^{N}$ are drawn i.i.d. from ¯$\bar{\mathbb{Q}}_{N}$.
To analyze the $\mbox{median}[X_{N}|\bar{\mathbb{Q}}_{N}]$, we first compute the mean and variance of $X_{N}$. It is easy to see that ${\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[X_{N}]=1$, while the variance of $X_{N}$ can be bounded as follows:
$\begin{array}[]{ll}\mbox{Var}_{\bar{\mathbb{Q}}_{N}}(X_{N})&=\displaystyle\frac{\mbox{Var}_{\bar{\mathbb{Q}}_{N}}\left(\sum_{i=1}^{N}\exp(\alpha\zeta_{i})\right)}{\left(N{\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]\right)^{2}}\\[12.91663pt] &=\displaystyle\frac{\mbox{Var}_{\bar{\mathbb{Q}}_{N}}\left(\exp(\alpha\zeta)\right)}{N\left({\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]\right)^{2}}\\[12.91663pt] &\leq\displaystyle\frac{2G\exp(\frac{4G}{C})}{N(C-2)},\end{array}$
where the second equality follows from the fact that $\{\zeta_{i}\}_{i=1}^{N}$ are i.i.d. The inequality follows since {¯$\{\bar{\mathbb{Q}}_{N}\}_{N=1}^{\infty}$ satisfy Assumption 2, thus Lemma 1 provides bounds for both $\mbox{Var}_{\bar{\mathbb{Q}}_{N}}\left(\exp(\alpha\zeta)\right)$ and ${\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]$, resulting in the bound $\frac{\text{Var}_{\bar{\mathbb{Q}}_{N}}\left(\exp(\alpha\zeta)\right)}{\left({\mathbb{E}}_{\bar{\mathbb{Q}}_{N}}[\exp(\alpha\zeta)]\right)^{2}}\leq\frac{2G\exp(\frac{4G}{C})}{C-2}$.
We next show that $\mbox{median}[X_{N}|\bar{\mathbb{Q}}_{N}]$ is bounded. To this end, consider the Chebyshev’s inequality:
¯${\mathbb{P}}_{\bar{\mathbb{Q}}_{N}}\left(\lvert X_{N}-1\rvert\geq k\sqrt{\mbox{Var}_{\bar{\mathbb{Q}}_{N}}(X_{N})}\right)\leq\frac{1}{k^{2}}.$
Substituting the bound for the $\mbox{Var}_{\bar{\mathbb{Q}}_{N}}(X_{N})$ and setting $k=2$, which implies an upper tail probability bound of 25%, results in
@@
We conclude from the above analysis that any sequence {¯$\{\bar{\mathbb{Q}}_{N}\}_{N=1}^{\infty}$ that is uniformly exponentially bounded as prescribed in Assumption 2 must be such that $\exp(-\alpha\delta_{N}(\bar{\mathbb{Q}}_{N}))=\mbox{median}[X_{N}|\bar{\mathbb{Q}}_{N}]$ converges to 1 almost surely. Since Assumption 2 imposes that such sequences are almost surely obtained, we must have that $\delta_{N}(\mathbb{Q}_{N})\rightarrow 0$ almost surely.
A.3 Proof of Proposition 3
Proof. We will show that $\rho_{\text{{SAA}}}$ converges to $\rho^{*}$ by demonstrating that $\rho_{\hat{\mathbb{P}}_{N}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))$ converges to $\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))$ for all ${{\boldsymbol{z}}}\in\mathcal{Z}$, i.e., a uniform rate of convergence. The proof is conducted in two steps. In the first step, we apply Jiang et al., (2020, Theorem 3.2) to show that the empirical exponential utility, ${\mathbb{E}}_{\hat{\mathbb{P}}_{N}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]$, converges uniformly to the true exponential utility, ${\mathbb{E}}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]$. In the second step, we leverage the properties of the logarithm function to extend this uniform convergence from the empirical exponential utility to the entropic risk measure.
To apply Jiang et al., (2020, Theorem 3.2), we need to verify several properties for the exponential utility minimization problem. It is easy to see that $(i)$ $\mathcal{Z}$ is compact by Assumption 3(A.1); $(ii)$ the exponential utility ${\mathbb{E}}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]$ is continuous in $\boldsymbol{z}$ and $(iii)$ the empirical distribution ^$\hat{{\mathbb{P}}}_{N}$ is constructed using $N$ i.i.d samples from ${\mathbb{P}}$; $(iv)$ Assumption 3(A.5) implies that $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ satisfies Assumption 1 for all ${{\boldsymbol{z}}}\in\mathcal{Z}$ so that Lemma 1 confirms that ${\mathbb{E}}[\left(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})\right)^{2}]<\infty$ for all ${{\boldsymbol{z}}}\in\mathcal{Z}$. Lastly, $(v)$ the following inequalities verify that $\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))$ is H-calm from above with Lipschitz modulus $\kappa({{\boldsymbol{\xi}}}):=\alpha L({{\boldsymbol{\xi}}})\exp(\alpha\bar{L}({{\boldsymbol{\xi}}}))$ and order one:
$\exp(\alpha\ell({{\boldsymbol{z}}}_{2},{{\boldsymbol{\xi}}}))-\exp(\alpha\ell({{\boldsymbol{z}}}_{1},{{\boldsymbol{\xi}}}))$ $\ \leq\left(\left.\partial_{{\boldsymbol{z}}}\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))\right|_{{{\boldsymbol{z}}}={{\boldsymbol{z}}}_{2}}\right)^{\top}\left({{\boldsymbol{z}}}_{2}-{{\boldsymbol{z}}}_{1}\right)$
$\ =\alpha\exp(\alpha\ell({{\boldsymbol{z}}}_{2},{{\boldsymbol{\xi}}}))\left(\partial_{{{\boldsymbol{z}}}}\ell({{\boldsymbol{z}}}_{2},{{\boldsymbol{\xi}}})\right)^{\top}\left({{\boldsymbol{z}}}_{2}-{{\boldsymbol{z}}}_{1}\right)$
$\ \leq\ \alpha\exp(\alpha\ell({{\boldsymbol{z}}}_{2},{{\boldsymbol{\xi}}}))\|\partial_{{\boldsymbol{z}}}\ell({{\boldsymbol{z}}}_{2},{{\boldsymbol{\xi}}})\|_{*}\|{{\boldsymbol{z}}}_{2}-{{\boldsymbol{z}}}_{1}\|$
$\ \leq\ \alpha\exp(\alpha\bar{L}({{\boldsymbol{\xi}}}))L({{\boldsymbol{\xi}}})\lVert{{\boldsymbol{z}}}_{2}-{{\boldsymbol{z}}}_{1}\rVert.$
The first inequality follows from the convexity of the exponential utility, where $\partial_{{\boldsymbol{z}}}$ denotes the subgradient operator with respect to ${{\boldsymbol{z}}}$, and the equality follows by applying the chain rule. The second inequality follows from the definition of the dual norm. Since by Assumption 3 $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is locally Lipschitz continuous with Lipschitz constant $L({{\boldsymbol{\xi}}})$, and the dual norm of the subgradient of $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is bounded by the Lipschitz constant $L({{\boldsymbol{\xi}}})$ (Shalev-Shwartz et al.,, 2012, Lemma 2.6), we obtain the final inequality. Finally, we can confirm that the second moment of $\kappa(\boldsymbol{\xi})$ is bounded. Namely, letting $10<\varsigma<(C/2)-$ 1, based on Hölder’s inequality, we have that:
${\mathbb{E}}[\kappa({{\boldsymbol{\xi}}})^{2}]$ $=\alpha^{2}{\mathbb{E}}[L({{\boldsymbol{\xi}}})^{2}\exp(2\alpha\bar{L}({{\boldsymbol{\xi}}}))]\leq\alpha^{2}{\mathbb{E}}[L({{\boldsymbol{\xi}}})^{2(1+1/\varsigma)}]^{\varsigma/(1+\varsigma)}{\mathbb{E}}[\exp(2(1+\varsigma)\alpha\bar{L}({{\boldsymbol{\xi}}}))]^{1/(1+\varsigma)}$
We can further show using Assumption 3(A.4) that :